OpenStudy (daisyblake11):

use the given values to evaluate (if possible) all six trig functions: cos (pi/2 - x) and cosx = 12/13

Directrix (directrix):

Is this the question as it appears in your text? Could you post a screen shot of the problem? Something seems to be missing.

Directrix (directrix):

^^ @Daisyblake11

OpenStudy (daisyblake11):

yes that's how it is

OpenStudy (daisyblake11):

one sec

Directrix (directrix):

>>cos (pi/2 - x) Is this complete?

OpenStudy (daisyblake11):

lol it isn't. my bad

Directrix (directrix):

cosx = 12/13 is partially helpful but we do not know whether we are in quadrant III or quadrant I. If you post a screen shot of the problem, I may be able to help. Otherwise, I cannot.

OpenStudy (daisyblake11):

it's cos(pi/2-x) = 5/13 and cos x = 12/13

Directrix (directrix):

cos (pi/2 - x) = sin x = 5/13 So far, we have: cos x = 12/13 sin x = 5/13 That leaves: sec x csc x tan x cot x

Directrix (directrix):

@Daisyblake11 Do you know the reciprocals of the trig functions?

OpenStudy (daisyblake11):

yes i do. thank yo so much

Directrix (directrix):

Good to know. What is this equal: sec x =

OpenStudy (daisyblake11):

13/12

Directrix (directrix):

csc x =

OpenStudy (daisyblake11):

13/5

Directrix (directrix):

tan x =

OpenStudy (daisyblake11):

sinx/cosx

Directrix (directrix):

Right. Substitute the values we already have for sin x and cos x into the tangent fraction and crank out the numerical value of the tangent.

OpenStudy (daisyblake11):

got it

OpenStudy (daisyblake11):

5/12

Directrix (directrix):

cot x =

OpenStudy (daisyblake11):

12/5

Directrix (directrix):

We have them all. Good work.

OpenStudy (daisyblake11):

thank you. good work :)

Directrix (directrix):

You are welcome.